Dickson, Inc., has a debt-equity ratio of 2.5. The firm’s weighted average cost of capital is 11 percent and its pretax cost of debt is 9 percent. The tax rate is 22 percent.

  

a. What is the company’s cost of equity capital? (Do not round intermediate calculations and enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.)
b. What is the company’s unlevered cost of equity capital? (Do not round intermediate calculations and enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.)
c. What would the company’s weighted average cost of capital be if the company's debt-equity ratio were .60 and 1.50? (Do not round intermediate calculations and enter your answers as a percent rounded to 2 decimal places, e.g., 32.16.)
Answer #1
D/A = D/(E+D)
D/A = 2.5/(1+2.5)
=0.7143
E/A = 1-D/A
=1-0.7143
=0.2857
a
WACC = Levered cost of equity*E/A+Cost of debt*(1-tax rate)*D/A
0.11= Levered cost of equity*0.2857+0.09*(1-0.22)*0.7143
Levered cost of equity =-1.16%
b
Levered cost of equity = Unlevered cost of equity+D/E*( Unlevered cost of equity-cost of debt)*(1-tax rate)
-0.0115852975495916 = Unlevered cost of equity+2.5*(Unlevered cost of equity-0.09)*(1-0.22)
Unlevered cost of equity = 5.56
C
Levered cost of equity = Unlevered cost of equity+D/E*( Unlevered cost of equity-cost of debt)*(1-tax rate)
Levered cost of equity = 5.56+0.6*(5.56-0.09)*(1-0.22)
Levered cost of equity = 8.12
C
Levered cost of equity = Unlevered cost of equity+D/E*( Unlevered cost of equity-cost of debt)*(1-tax rate)
Levered cost of equity = 5.56+1.5*(5.56-0.09)*(1-0.22)
Levered cost of equity = 11.96